User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Patrick

Next meeting chair: Katie

Attended: Jay, Eloise, Jamie, Lim, Prad, Mas, Simin, Warren


General Issues

  • Don't submit checklists for Ho - still broken.
  • AUM observing notes should be pretty good
    • sx time test script is necessary
  • Ho had problem with LO being unlocked for last ~3 experiments. Error was suppressed in logs. rxmon reports whether LO is locked/unlocked, but was not picked up in the logs.
    • Make sure that when suppressing cryo temperatures, we don't suppress LO locking status, as this is a critical error.

Handover Issues

  • AUM018
    • Hb:
      • Incorrect procedure file used
    • Ke:
      • dbbcke gave errors during configuring, had to restart control software (same version though)
      • Issues with which polarisations were recording
        • Ongoing problems with dbbc2, internal Fila10g is not staying configured. Not sure of origin yet.
      • Fila10g crashed
  • RD2007
    • Ho:
      • Communication issues with mk4 rack (?)
/home/www/auscope/opswiki/data/pages/operations/agenda2020_08_27.txt · Last modified: 2020/08/27 03:16 by Patrick Yates